OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions. |s + 12} = 15

OpenStudy (lgbasallote):

solve for s in s+ 12 = 15 and s + 12 = -15 does that help?

OpenStudy (anonymous):

It says the answer is -27.3 ... and I made a typo. * |s + 12 | = 15

OpenStudy (lgbasallote):

i still say the same thing